OpenStudy (anonymous):

A bean of unknown elementary particles travels at a speed of 2 x 10^8 m/s. Their average lifetime in the beam is measured to be 1.6 x 10^-8 s. Calculate their average lifetime when at rest.

OpenStudy (johnweldon1993):

hmm...it's definitely been awhile since I've done these...I believe you start with the equation \[t = \frac{ t_o }{ \sqrt{1-\frac{ v^2 }{ c^2 }} }\] We want t_o so rearranging we have \[t_o = t \sqrt{1-\frac{ v^2 }{ c^2 }}\] You know v = speed it's traveling at t = average lifetime c = speed of light so replace v with 2 x 10^8 replace t with 1.6 x 10^-8 and replace c with 3.0 x 10^8 to get \[(1.6 * 10^{-8})\sqrt{1-\frac{ (2.0 * 10^8) }{ (3.0 * 10^8 )}^2}\] *and yes that ^2 is outside of that whole fraction...it should be [(2.0 * 10^8)/(3.0 * 10^8)] all ^2 So can you solve that?

OpenStudy (anonymous):

thank you

OpenStudy (johnweldon1993):

no problem...I believe that is correct...better you double check my work...as stated it HAS been a while since doing those types of problems

OpenStudy (anonymous):

can you help with another problem